### v2

# Social Bias Probing: Fairness Benchmarking for Language Models

Abstract
Language models have been shown to encode a variety of social biases, which carries the risk of downstream harms. While the impact of these biases has been recognized, prior methods for bias evaluation have been limited to binary association tests on small datasets, offering a constrained view of the nature of social biases within language models. In this paper, we propose a novel framework for probing language models for social biases. We collect a probing dataset to analyze language models’ general associations, as well as along the axes of social categories, identities, and stereotypes. To this end, we leverage a novel perplexity-based fairness score. We curate a large-scale benchmarking dataset addressing the limitations of existing fairness collections, expanding to a variety of different identities and stereotypes. When comparing our methodology with prior work, we demonstrate that biases within language models are more nuanced than previously acknowledged, suggesting that the scope of biases LMs encode is broader than previously understood. Moreover, we expose how identities expressing different religions lead to the most pronounced disparate treatments across all models.
1 Introduction
The unparalleled ability of language models to generalize from vast corpora is tinged by an inherent reinforcement of social biases. These biases are not merely encoded within language models’ representations but are also perpetuated to downstream tasks (Blodgett et al., 2021; Stańczak and Augenstein, 2021), where they can manifest in an uneven treatment of different demographic groups – a challenge documented across various studies (Rudinger et al., 2018; Stanovsky et al., 2019; Kiritchenko and Mohammad, 2018; Venkit et al., 2022).
Figure 1: Workflow of Social Bias Probing Framework.
A direct analysis of biases encoded within language models allows to pinpoint the problem at its source, potentially obviating the need for addressing it for every application (Nangia et al., 2020). Therefore, a number of studies have attempted to evaluate social biases within language models (Nangia et al., 2020; Nadeem et al., 2021; Stańczak et al., 2023; Nozza et al., 2022a). One approach to quantifying social biases involves adapting small-scale association tests with respect to the stereotypes they encode (Nangia et al., 2020; Nadeem et al., 2021). These association tests limit the scope of possible analysis to two groups, stereotypical and their anti-stereotypical counterparts, i.e., the identities that “embody” the stereotype and the identities that violate it, often connotated with different/opposite beliefs. This binary approach, which assumes a singular “ground truth” with respect to a stereotypical statement, has restricted the depth of the analysis and simplified the complexity of social identities and their associated stereotypes. The nuanced nature of social biases within language models has thus been largely unexplored.
To overcome these limitations, we introduce a novel probing framework, as outlined in Figure 1. The input of our approach consists of a dataset gathering stereotypes and a set of identities belonging to different social categories: gender, religion, disability, and nationality. First, we combine stereotypes and identities resulting in our probing dataset. Secondly, we assess social biases across five language modeling architectures in English. We propose perplexity (Jelinek et al., 1977), a measure of a language model’s uncertainty, as a proxy for bias. By evaluating how a language model’s perplexity varies when presented with probes that contain identities belonging to different social categories, we can infer which identities are considered the most likely. Using the perplexity-based fairness score, we conduct a three-dimensional analysis: by social category, identity, and stereotype for each of the considered language models. In summary, the contributions of this work are:
  • We conceptually facilitate fairness benchmarking across multiple identities going beyond the binary approach of a stereotypical and an anti-stereotypical identity.
  • We deliver SoFa (Social Fairness), a benchmark resource to conduct fairness probing addressing drawbacks and limitations of existing fairness datasets, expanding to a variety of different identities and stereotypes. [^1] SoFa will be made available upon paper acceptance.
  • We propose a perplexity-based fairness score to measure language models’ associations with various identities.
  • We study social biases encoded within five different language modeling architectures along the axes of social categories, identities, and stereotypes.
A comparative analysis with the popular benchmarks CrowS-Pairs Nangia et al. (2020) and StereoSet Nadeem et al. (2021) reveals marked differences in the overall fairness ranking of the models, suggesting that the scope of biases LMs encode is broader than previously understood. Moreover, we expose how identities expressing religions lead to the most pronounced disparate treatments across all models, while the different nationalities appear to induce the least variation compared to the other examined categories, namely gender and disability.
2 Related Work
Presenting a recent framing on the fairness of language models, Navigli et al. (2023) define social bias [^2] The term social characterizes bias in relation to the risks and impacts on demographic groups, distinguishing it from other forms of bias, e.g. the statistical one. as the manifestation of “prejudices, stereotypes, and discriminatory attitudes against certain groups of people” through language. Social biases are featured in training datasets and propagated in downstream NLP applications, where it becomes evident when the model exhibits significant errors in classification settings for specific minorities or generates harmful content when prompted with sensitive identities.
Fairness Scores and Datasets
Delobelle et al. (2022) report various bias metrics for pre-trained LMs. The HONEST score, proposed by Nozza et al. (2021), is computed on template-based sentences created to measure the hurtfulness of LMs’ completions within the task of masked language modeling.
Recent work Blodgett et al. (2021) has pointed out relevant concerns regarding stereotype framing and data reliability of benchmark collections explicitly designed to analyze biases in language models, such as CrowS-Pairs Nangia et al. (2020) and StereoSet Nadeem et al. (2021). The scores proposed in the contributions presenting the datasets are highly dependent on the form of the resource they propose and therefore they are hardly generalizable to other datasets to conduct a more general comparative analysis. Specifically, Nangia et al. (2020) leverage pseudo-log likelihood Salazar et al. (2020) based scoring. The score assesses the likelihood of the unaltered tokens based on the modified tokens’ presence. It quantifies the proportion of instances where the LM favors the stereotypical sample (or, vice versa, the anti-stereotypical one). The stereotype score proposed by Nadeem et al. (2021) differs from the former as it allows bias assessment on both masked and autoregressive language models, whereas CrowS-Pairs is limited to the masked ones. Another significant constraint of both datasets is the establishment of a bias score threshold at $50\%$ (Pikuliak et al., 2023). It implies that a model displaying a preference for stereotypical associations more than $50\%$ of the time is considered biased, and vice versa. This threshold implies that a model falling below it may be deemed acceptable or, in other words, unbiased. Further, these datasets exhibit limitations regarding their focus and coverage of identity diversity and the number of stereotypes. This stems from their reliance on the binary comparison between two rigid alternatives — the stereotypical and anti-stereotypical associations — which fails to capture the phenomenon’s complexity. Indeed, they do not account for how the model behaves in the presence of other plausible contexts, and these potential associations need scrutiny for low probability generation by the model, as they can be harmful regardless of the specific target. Therefore, bias measurements using these resources could lead to unrealistic and inaccurate fairness evaluations.
Given the constraints of current solutions, our work introduces a dataset that encompasses a wider range of identities and stereotypes. The contribution relies on a novel framework for probing language models for social biases. To address the limitations identified in the literature review, we design a perplexity-based ranking that produces a more nuanced evaluation of fairness. Regarding adopting of perplexity as a fairness measure, we recognize similarities in the insights leveraged by All Unmasked Likelihood, a novel bias evaluation measure for masked LMs proposed by Kaneko and Bollegala (2022). Compared to AUL and other similar evaluations Nangia et al. (2020); Nadeem et al. (2021), (i) our evaluation approach does not assume but rather rejects the existence of a stereotypical and an anti-stereotypical phrase as framing to measure what is biased and what is not; (ii) we do not differ between modified and unmodified tokens; (iii) we contest the need for a threshold of $50$ considered as a neutral, desirable value; (iv) certain scores are specifically suitable for masked LMs, while we focus our analysis and our metric for autoregressive causal LMs.
3 Social Bias Probing Framework
The proposed Social Bias Probing framework serves as a technique for fine-grained fairness benchmarking of LMs. Contrary to the existing fairness assessments, which rely on a dichotomous framework of stereotypical vs. anti-stereotypical associations, our methodology expands beyond this oversimplified binary categorization. Ultimately, our approach enables the comprehensive evaluation of language models by incorporating a diverse array of identities, thus providing a more realistic and rigorous audit of fairness within these systems.
Figure 1 presents our workflow. We first collect a set of stereotypes and identities leveraging the Social Bias Inference Corpus (SBIC; Sap et al. 2020) and an identity lexicon curated by Czarnowska et al. (2021). At this stage, we develop the new SoFa (Social Fairness) dataset, which encompasses all probes — identity–stereotype combinations (Section 3.1). The final phase of our workflow involves evaluating language models by employing our proposed perplexity-based fairness measure in response to the constructed probes (Section 3.2).
3.1 Probing Dataset Generation
Our approach requires a set of identities from diverse social and demographic groups, alongside an inventory of stereotypes.
Stereotypes
We derive stereotypes from the list of implied statements in SBIC, a corpus that collects social media posts having harmful biased implications. The posts are sourced from previously published collections that include English content from Reddit and Twitter, for a total of $44,000$ instances. Additionally, the authors draw from two “Hate Sites”, namely Gab and Stormfront. Annotators were asked to label the texts based on a conceptual framework designed to represent implicit biases and offensiveness. [^3] We refer to the dataset for an in-depth description (https://maartensap.com/social-bias-frames/index.html).
We emphasize that SBIC serves as an exemplary instantiation of our framework. Our methodology can be applied more broadly to any dataset containing stereotypes directed towards specific categories. As not all instances of the original dataset have an annotation regarding the stereotype implied by the social media comment, we filter it to isolate abusive samples having a stereotype annotated. Since certain stereotypes contain the targeted identity, whereas our goal is creating multiple control probes with different identities, we remove the subjects from the stereotypes, performing a preprocessing to standardize the format of statements (details are documented in Appendix A). Finally, we discard stereotypes with high perplexity scores to remove unlikely instances [^4] Other works in the literature also adopt this procedure, e.g., Barikeri et al. (2021).. Since our evaluation is based on identifying and quantifying perplexity peaks (both low and high), keeping statements in the dataset that are unlikely according to the models would offset the results. The filtering is based on a threshold, specifically averaging perplexity scores from each model and creating a histogram to retain only stereotypes in selected bins exhibiting the lowest scores (Fig. 6).
Identities
While we could have directly used the identities provided in the SBIC dataset, we chose not to due to their unsuitability from frequent repetitions and varying expressions influenced by individual annotators’ styles. To unify the set of analyzed identities, we deploy the lexicon created by Czarnowska et al. (2021): Table 5 in Appendix reports a sample for each category. [^5] The complete list of identities is available at https://github.com/amazon-science/generalized-fairness-metrics/tree/main/terms/identity_terms. We map the SBIC dataset target group categories to the identities available in the lexicon (Table 3). Specifically, the categories are: gender, race, culture, disabilities, victim, social, and body. We first define and rename the culture category to include religions and broaden the scope of the race category to encompass nationalities. We then link the categories in the SBIC dataset to those present in the lexicon as follows: gender identities are extracted from the lexicon’s genders and sexual orientations, nationality identities are derived from race and country entries, religion utilizes terms from the religion category, and disabilities are drawn from the disability category. This mapping results in excluding the broader SBIC categories, i.e., victim, social, and body due to the difficulty in aligning the identities with the lexicon categories, and disproving the assumption of invariance in the related statements. [^6] This choice is motivated by the fact that the stereotypes under these categories are often specific to a particular identity; for example, they might have referenced body parts belonging to one gender and not another. The assignment of an identity to a specific category is inherited from the categorizations of the resources adopted. Recognizing that these framings inevitably simplify the complex real-world nuances is crucial.
SoFa
To obtain the final probing dataset, we remove duplicated statements and apply lower-case. Finally, each target is concatenated to each statement with respect to their category, creating dataset instances that differ only for the target (Table 6). In Table 3, we report the coverage statistics regarding targeted categories and identities.
3.2 Fairness Measure
Measure
We propose perplexity (PPL; Jelinek et al. 1977) as a means of intrinsic evaluation of fairness in language models. PPL is defined as the exponentiated average negative log-likelihood of a sequence. More formally, let $X=(x_{0},x_{1},\dots,x_{t})$ be a tokenized sequence, then the perplexity of the sequence is
$PPL(X)=\exp\left\{-\frac{1}{t}\sum_{d}^{t}\log_{e}p_{\theta}(x_{d}\mid x_{<d})\right\}$
where $\log p_{\theta}(x_{d}\mid x_{<d})$ is the log-likelihood of the $d$th token conditioned on the proceeding tokens given a model parametrized with $\theta$.
Our metric leverages PPL to quantify the propensity of a model to produce a given input sentence: a high PPL value suggests that the model deems the input improbable for generation. We identify bias manifestations when a model exhibits low PPL values for statements that contain stereotypes, thus indicating a higher probability of their generation. The purpose of this metric, and more generally our framework, is to provide a fine-grained summary of models’ behaviors from an invariance fairness perspective, i.e., the same statement referring to different demographic groups should not cause a significant change in model behavior, or, in more general terms, individuals from different demographic groups should be treated equally.
[…]
${PPL^{\star m}_{(i+s)}}=\frac{PPL^{m}_{(i+s)}}{PPL^{m}_{(i)}}$ (1)
Since the identities $i$ are characterized by their own PPL scores, we normalize the PPL of the probe with the PPL of the identity, addressing the risk that certain identities might yield higher PPL scores because they are considered unlikely.
We highlight that the PPL’s scale across different models can significantly differ based on the training data. Consequently, the raw scores do not allow direct comparisons. We facilitate the comparison of the PPL values of model $m_{1}$ and model $m_{2}$ for a given combination of identity and a stereotype:
$PPL^{\star m_{1}}_{(i+s)}\equiv k\cdot{PPL^{\star m_{2}}_{(i+s)}}$ (2)
$\log_{10}({PPL^{\star m_{1}}_{(i+s)}})$ $\equiv\log_{10}(k\cdot{PPL^{\star m_{2}}_{(i+s)}})$ (3)
$\sigma^{2}(\log_{10}({PPL^{\star m_{1}}_{P_{c,s}}}))$ $=\sigma^{2}(\log_{10}(k)+\log_{10}({PPL^{\star m_{2}}_{P_{c,s}}}))$
$=\sigma^{2}(\log_{10}{PPL^{\star m_{2}}_{P_{c,s}}})$ (4)
In Eq. 2, $k$ is a constant and represents the factor that quantifies the scale of the scores emitted by the model: in fact, different models emit scores having different scales and, therefore, as already mentioned, are not directly comparable. Importantly, each model has its own $k$ [^7] The constant $k$ is not calculated; it is only formally described. The assumption of the existence of this constant $k$ allows us to compare perplexity values through the steps outlined., but because it is a constant, it does not depend on the input text sequence but solely on the model $m$ in question. In Eq. 3, we use the base-$10$ logarithm of the PPL values generated by each model to analyze more tractable numbers since the range of PPL is $[0,\inf)$. For the purpose of calculating variance across the probes $P_{c,s}$ (Eq. 3.2), which is the main investigation conducted in our dataset, $k$ plays no role and does not influence the result. Consequently, we can compare different PPLs from models that have been transformed in this manner.
Lastly, we define Delta Disparity Score (DDS) as the magnitude of the difference between the highest and lowest PPL score as a signal for a model’s bias with respect to a specific stereotype. DDS is computed separately for each stereotype $s$ belonging to category $c$, or, in other words, on the set of probes created from the stereotype $s$.
$DDS_{P_{c,s}}=\max_{P_{c,s}}(\log({PPL^{\star m}_{(i+s)}}))$ (5)
$-\min_{P_{c,s}}(\log({PPL^{\star m}_{(i+s)}}))$
Evaluation
We conduct the following types of evaluation: intra-identities, intra-stereotypes, intra-categories, and calculate a global fairness score. At a fine-grained level, we identify the most associated sensitive identity intra-$i$, i.e., for each stereotype $s$ within each category $c$. This involves associating the $i$ achieving the lowest (top-$1$) $\log({PPL^{\star m}_{(i+s)}})$ as reported in Equation 3, PPL from now on for the sake of brevity. Additionally, we delve into the analysis of stereotypes themselves (intra-$s$), exploring DDS as defined in Equation 5 between the maximum and minimum PPLs obtained for the set of probes generated from $s$ (again, for each $c$, for each $s$ within $c$). This comparison allows us to pinpoint the strongest stereotypes within each category (in the sense of the ones causing the lowest disparity w.r.t. DDS), shedding light on the shared stereotypes across identities. Extending our exploration to the intra-category level, we aggregate and count findings from the intra-identities and stereotypes settings. At a broader level, our goal is to uncover, for each sensitive category, the top-$k$ strongest (low PPL) identities and stereotypes within that category. The findings resulting from the various settings are first investigated separately for each model $m$. In the subsequent analysis, we delve into the overlap among the top-$k$ identities and stereotypes, spanning both within and across model families and scales.
To obtain a global fairness score for each $m$, for each $c$ and $s$ we compute the variance as formalized in Section 3.2 occurring among the probes of $s$, and average it by the number of $s$ belonging to $c$. Having computed the variance for $c$, we perform a simple average to obtain the final number. This aggregated number finally allows us to compare the behavior of the various models on the dataset and to rank the models according to variance: models reporting a higher variance are more unfair. We reference this measure as SoFa score.
4 Experiments and Results
In this work, we benchmark five auto-regressive causal language models using our framework: BLOOM (Scao et al., 2022), GPT2 (Radford et al., 2019), XLNET (Yang et al., 2019), BART (Lewis et al., 2020), and LLAMA2 (Touvron et al., 2023). We opt for models accessible through the Hugging Face Transformers library Wolf et al. (2020), among the most recent, popular, and demonstrating state-of-the-art performance across various NLP tasks. Our selection process also involved considering language models audited by other fairness benchmark datasets, specifically StereoSet Nadeem et al. (2021) and CrowS-Pairs Nangia et al. (2020), to enable direct comparison. In Table 4, we list the selected language models [^8] Due to computational constraints, we did not include more different models. Future work could include additional models: we would expect the findings to be similar to those we have already documented.: for each, we examine two scales with respect to the number of parameters. The PPL is computed at the token level through Hugging Face’s evaluate library. [^9] https://huggingface.co/spaces/evaluate-metric/perplexity
4.1 Benchmarks
We compare our framework against two other popular fairness benchmarking frameworks previously introduced in Section 2. StereoSet (Nadeem et al., 2021): To assess the bias in a language model, the model is scored using likelihood-based scoring of the stereotypical or anti-stereotypical association in each example. The percentage of examples where the model favors the stereotypical association over the anti-stereotypical one is calculated as the model’s stereotype score. CrowS-Pairs (Nangia et al., 2020): The bias of a language model is assessed by evaluating how often it prefers the stereotypical sentence over the anti-stereotypical one in each pair using pseudo-likelihood-based scoring. Compared to these evaluation methodologies, our metric does not impose an artificial threshold. Our perplexity-based approach overcomes the limitation of a fixed threshold, such as $\theta=50\%$, i.e., if a model prefers stereotypical associations exceeding $\theta$, it is deemed unfair. We can therefore investigate the behavior of models in a more nuanced and less constrained manner. Our multifaceted approach allows us to gain insights into the complex relationships between identities and stereotypes across categories and models.
4.2 Results
Models Datasets
SoFa ($1.633.477$) StereoSet ($4,229$) CrowS-Pairs ($1,508$)
Family Size Rank $\downarrow$ Score $\downarrow$ Rank $\downarrow$ Score $\downarrow$ Rank $\downarrow$ Score $\downarrow$
BLOOM 560m 1 $2.325$ 6 $7.925$ $58.91$
3b 9 $0.33$ 4 $61.11$ 4 $61.71$
GPT2 base 7 $0.361$ 5 $0.426$ $58.45$
medium 8 $0.35$ 3 $62.91$ 3 $63.26$
XLNET base 4 $0.795$ 8 $52.20$ 7 $49.84$
large 2 $1.422$ 7 $53.88$ 8 $48.76$
BART base 10 $\mathbf{0.072}$ 10 $\mathbf{47.82}$ 10 $\mathbf{39.69}$
large 3 $0.978$ 9 $51.04$ 9 $44.11$
LLAMA2 7b 6 $0.374$ 2 $63.36$ 2 $70$
13b 5 $0.387$ 1 $64.81$ 1 $71.32$
Table 1: Results on SoFa and the two previous fairness benchmarks, StereoSet and CrowS-Pairs. The ranking ranges from 1 (LM most biased) to 10 (LM least biased $\downarrow$); for each of the scores, the best value in bold is the lowest $\downarrow$, connoting the least biased model. We note the number of instances in each dataset next to their names.
Model Category normal-$\downarrow$
Family Size Relig. Gend. Dis. Nat.
BLOOM 560m $3.216$ $2.903$ $1.889$ $1.292$
3b $0.376$ $0.483$ $0.301$ $0.162$
GPT2 base $0.826$ $0.340$ $0.161$ $0.116$
medium $0.839$ $0.304$ $0.164$ $0.091$
XLNET base $0.929$ $0.803$ $0.846$ $0.601$
large $2.044$ $1.080$ $1.554$ $1.012$
BART base $\mathbf{0.031}$ $\mathbf{0.080}$ $\mathbf{0.107}$ $\mathbf{0.071}$
large $1.762$ $1.124$ $0.582$ $0.442$
LLAMA2 7b $0.612$ $0.422$ $0.324$ $0.138$
13b $0.740$ $0.372$ $0.312$ $0.123$
Table 2: SoFa score by category (best value in bold).
Global fairness score evaluation
In Table 1, we report the results of our comparative analysis using the previously introduced benchmarks, StereoSet and CrowS-Pairs. [^10] In order to obtain the results, we used the implementation from https://github.com/McGill-NLP/bias-bench by Meade et al. (2022). The reported scores are based on the respective datasets, each of which incorporates different insights and interpretations of biases. Since the measures of the three fairness benchmarks are not directly comparable, we include a ranking column, ranging from 1 (most biased) to 10 (least biased). In fact, the ranking setting in the two other fairness benchmarks reports a percentage, as described in Section 2, whereas our score represents the average of the variances obtained per probe, as detailed in Section 3.2. Given that few values stand below $50$, a value considered neutral, ideal, according to StereoSet and CrowS-Pairs, we intuitively choose to interpret the best score as the lowest, consistent with SoFa’s assessment, and choose to consider a model slightly skewed toward the antistereotypical association as best rather than the other way around.
Through the ranking, we observe an exact agreement between StereoSet and CrowS-Pairs on the model order for the first four positions. Discrepancies are registered from the fifth to the eighth positions, which are swapped. For the last two, instead, they are again identical for BART). The score magnitudes are also similar for the positions in the middle (from the third to the sixth); in comparison, the other couples exhibit a more pronounced difference especially for the first and last ranks (LLAMA2 and BART). In contrast, the ranking provided by SoFa reveals differences in the overall fairness ranking of the models, suggesting that the scope of biases language models encode is broader than previously understood. [^11] We used Kendall’s Tau test to quantify the similarity of rankings: StereoSet and CrowS-Pairs achieve a value close to $1$ ($0.911$), indicating strong agreement, while both benchmarks compared to SoFa reach $-0.022$, a value that confirms the already recognized disagreement. A first distinction with the two prior fairness benchmarks is evident: unlike them, who regard both dimensions of LLAMA2 as the most biased, our assessment places the pair of models in the middle of the ranking. Moreover, unlike StereoSet and CrowS-Pairs where for three out of five models (namely XLNET, BART and LLAMA2) contiguous positions are occupied by models belonging to the same family, the rank emerging from our dataset exhibits a contrasting pattern, except for GPT2 and LLAMA2. For three out of five models, the larger variant exhibits more bias, corroborating the findings of previous research (Bender et al., 2021). This pattern is not mirrored by BLOOM and GPT2. According to SoFa, BLOOM-560m emerges as the model with the highest variance. Notably, and similarly to BART, the two sizes of the model stand at opposite poles of the ranking (1-9 and 10-3), showing the impact of scale. After it, XLNET-large holds the second position. Indeed, prior work identifies XLNET to be highly biased compared to other language model architectures (Stańczak et al., 2023). They are followed (at a distance) by BART-large. Conversely, BART-base attains the lowest score, securing the tenth position. This aligns with the rankings provided by StereoSet and CrowS-Pairs, although the disparities with the scores from other models are less pronounced in these benchmarks compared to our setting. The differences between our results and those from the two other fairness benchmarks could stem from the larger scope and size of our dataset.
Intra-categories evaluation
In the following, we analyze the results obtained on the SoFa dataset broken down by category [^12] Since the categories in SoFa are different and do not correspond to the two competitor datasets, in the absence of a one-to-one mapping, we do not report this disaggregated result for StereoSet and CrowS-Pairs., detailed in Table 2. In Figure 5 in the Appendix, we report the score distribution across categories and LMs through violin plots. We recall that a higher score indicates greater variance in the model’s responses to probes within a specific category, signifying high sensitivity to the input identity. For the two scales of BLOOM, we notice scores that are far apart when comparing the pairs of results obtained by category: this behavior is recorded by the previous overall ranking, which, as we have reported in Table 1, places these two models at opposite poles of the scale. In the case of GPT2, we observe a notably higher score in the religion category, while other categories exhibit similar magnitudes. Regarding XLNET-base, both religion and disability achieve the highest similar values. Compared, nationality diverge considerably less. Similarly to the base version of the model, XLNET-large demonstrates significantly stronger variance for religion and disability when contrasted with scores for others, particularly nationality, which records the lowest value, therefore indicating minor variability concerning those identities. Similarly, for BART-large and the two scales of LLAMA2, religion consistently emerges as the category causing the most distinct behavior compared to other identities. Therefore, across all models except for BLOOM-3b, religion consistently stands out as the category leading to the most pronounced disparate treatment, while nationality often attains the lowest value.
Figure 2: Stereotypes with lowest DDS according to Equation 5, per category.
Intra-identities evaluation
In Table 7, we report a more qualitative result, i.e., the identities that, in combination with the stereotypes, obtain the lowest PPL score: intuitively, the probes that each model is more likely to generate for the set of stereotypes afferent to that category. We highlight that the four categories of SoFa are derived by combining categories of both SBIC, the dataset used as a source of stereotypes, and the lexicon used for identities. Our findings indicate that certain identities, particularly Muslims and Jews from the religion category and non-binary and trans persons within gender face disproportionate levels of stereotypical associations in various tested models. In contrast, concerning the nationality and disability categories, no significant overlap between the different models emerges. A contributing factor might in the varying sizes of the identity sets derived from the lexicon used for constructing the probes, as detailed in Table 3.
Intra-stereotypes evaluation
We display in Figure 2 the top one stereotype reaching the lowest DDS, as per Equation 5, essentially reporting the most prevalent shared stereotypes across identities within each category (Table 8 in the Appendix presents a more-in depth overview documenting the top three stereotypes for each model). In the religion category, the most frequently occurring stereotype revolves around immoral acts and beliefs or judgments of repulsion. For the gender category, mentions of stereotypical behaviors and sexual violence are consistently echoed across models, while in the nationality category, references span to the lack of employment, physical violence (both endured and performed), crimes, and various other offenses. Stereotypes associated with disability encompass judgments related to appearance, physical incapacity, and other detrimental judgments.
5 Conclusion
This study proposes a perplexity-based scoring and a novel probing framework to capture social biases by auditing language models on an original fairness benchmark. A comparative analysis with the popular benchmarks CrowS-Pairs Nangia et al. (2020) and StereoSet Nadeem et al. (2021) reveals marked differences in the overall fairness ranking of the models, suggesting that the scope of biases LMs encode is broader than previously understood. Moreover, our findings suggest that certain identities, particularly Muslims and Jews from the religion category and non-binary and trans persons within gender, face disproportionate levels of stereotypical associations in various tested models. Further, we expose how identities expressing religions lead to the most pronounced disparate treatments across all models, while the different nationalities appear to induce the least variation compared to the other examined categories, namely, gender and disability. Consequently, we stress the need for a broader holistic bias analysis and mitigation.
Limitations
Our framework’s reliance on the fairness invariance assumption is a critical limitation, particularly since sensitive real-world statements often acquire a different connotation based on a certain gender or nationality, due to historical or social context. Therefore, associating certain identities with specific statements may not be a result of a harmful stereotype, but rather a portrayal of a realistic scenario. Moreover, relying on a fully automated pipeline to generate the probes could introduce inaccuracies, both at the level of grammatical plausibility (syntactic errors) and semantic relevance (e.g., discarding neutral statements that do not contain stereotypical beliefs): conducting a human evaluation of a portion of the synthetically generated text will be pursued.
Another simplification, as highlighted in Blodgett et al. (2021), arises from “treating pairs equally”. Treating all probes with equal weight and severity is a limitation of this work.
Generating statements synthetically, for example, by relying on lexica, carries the advantage of artificially creating instances of rare, unexplored phenomena. Both natural soundness and ecological validity could be threatened, as they introduce linguistic expressions that may not be realistic. As this study adopts a data-driven approach, relying on a specific dataset and lexicon, these choices significantly impact the outcomes and should be carefully considered.
Concerning the use of perplexity as a proxy for bias, in Section 3.2, we explained the transformations designed to render the scores of the different models more comparable, mainly to calculate variance. In this study, we compared our metric with the bias measures proposed by StereoSet and CrowS-Pairs; we refer to further comparisons with others coined in the literature for future investigations.
While our framework could be extended to languages beyond English, our experiments focus on the English language due to the limited availability of datasets for other languages having stereotypes annotated. We strongly encourage the development of multilingual datasets for probing bias in language models, as in Nozza et al. (2022b); Touileb and Nozza (2022); Martinková et al. (2023).
For future research, we aim to diversify the dataset by incorporating stereotypes beyond the scope of a U.S.-centric perspective as included in the source dataset for the stereotypes, SBIC. Additionally, we highlight the need for analysis of biases along more than one axis. We will explore and evaluate intersectional probes that combine identities across different categories. Lastly, considering that fairness measures investigated at the pre-training level may not necessarily align with the harms manifested in downstream applications Pikuliak et al. (2023), it is recommended to include an extrinsic evaluation, as suggested by prior work Mei et al. (2023); Hung et al. (2023).
Ethical Considerations
Our benchmark is highly reliable on the set of stereotypes and identities included in the probing dataset. We opted to use the list of identities from Czarnowska et al. (2021). However, the identities included encompass a range of perspectives that the lexicon in use may not fully capture. The stereotypes we adopt are derived from the collection by Sap et al. (2020), which aggregated potentially biased content from a variety of online platforms such as Reddit, Twitter, and specific hate sites. These platforms tend to be frequented by certain demographics. These forums, despite having a broader demographic than traditional media sources such as newsrooms, Wikipedia editors, or book authors (Wagner et al., 2015), predominantly reflect the biases and perspectives of white men from Western societies.
In light of these considerations, following Attanasio et al. (2022), we advocate for the responsible use of benchmarking suites. Our benchmark is intended to be a starting point, and we recommend its application in conjunction with human-led evaluations. Users are encouraged to further develop and refine our probing dataset to enhance its inclusivity in terms of identities, stereotypes, and models included.
Acknowledgements
This research was co-funded by Independent Research Fund Denmark under grant agreement number 9130-00092B, and supported by the Pioneer Centre for AI, DNRF grant number P1. The work has also been supported by the European Community under the Horizon 2020 programme: G.A. 871042 SoBigData++, ERC-2018-ADG G.A. 834756 XAI, G.A. 952215 TAILOR, and the NextGenerationEU programme under the funding schemes PNRR-PE-AI scheme (M4C2, investment 1.3, line on AI) FAIR (Future Artificial Intelligence Research).
Appendix A Preprocessing
To standardize the format of statements, we devise a rule-based dependency parsing from a manual check of approximately $250$ stereotypes. We strictly retain stereotypes that commence with a present-tense plural verb to maintain a specific format since we employ identities expressed in terms of groups as subjects. Singular verbs are modified to plural for consistency using the inflect package. [^13] https://pypi.org/project/inflect/ We exclude statements that already specify a target, refer to specific recurring historical events, lack verbs, contain only gerunds, expect no subject, discuss terminological issues, or describe offenses and jokes rather than stereotypes.
We also preprocess the collected identities from the lexicon to ensure consistency regarding part-of-speech (PoS) and number (singular vs. plural). Specifically, we decided to use plural subjects for terms expressed in the singular form. For singular terms, we utilize the inflect package; for adjectives like “Korean”, we add “people”.
Appendix B Data Statement
We provide a data statement of SoFa, as proposed by Bender and Friedman (2018). We also report in Table 6 the dataset structure.
Curation Rationale
SoFa dataset consists of combined stereotypes and identities. The stereotypes are sourced from the SBIC dataset, and we refer the reader to Sap et al. (2020) for an in-depth description of the data collection process. For insights into the identities incorporated within SoFa, see Czarnowska et al. (2021).
Language Variety
en-US. Predominantly US English, as written in comments on Reddit, Twitter, and hate communities included in the SBIC dataset.
Author and Annotator Demographics
We inherit the demographics of the authors of the stereotypes and the annotations from Sap et al. (2020).
Text Characteristics
The analyzed stereotypes are extracted from the SBIC dataset. This dataset includes annotated English Reddit posts, specifically three intentionally offensive subReddits and a corpus of potential microaggressions from Breitfeller et al. (2019), and posts from three existing English Twitter datasets annotated for toxic or abusive language (Founta et al., 2018; Waseem and Hovy, 2016; Davidson et al., 2017). Finally, SBIC includes posts from known English hate communities: Stormfront (de Gibert et al., 2018) and Gab  [^14] https://files.pushshift.io/gab/GABPOSTS_CORPUS.xz which are both documented white-supremacist and neo-nazi communities and two English subreddits that were banned for inciting violence against women (r/Incels and r/MensRights).
Provenance
Data statements were not provided with SBIC, the underlying source of the stereotypes.
Appendix C Supplementary Material
In Table 3, we report the coverage statistics regarding targeted categories and identities. In Table 4, we list the selected language models: for each, we examine two scales with respect to the number of parameters. In Table 5, we report example identities for each category of the SoFa dataset. We deploy the lexicon created by Czarnowska et al. (2021): the complete list is available at https://github.com/amazon-science/generalized-fairness-metrics/tree/main/terms/identity_terms.
In Figure 3, we provide an illustrative example to convey the fundamental insight of our perplexity-based scoring. In Figure 4, we display the SoFa score by category (numbers in Table 2).
In Table 7, we report the identities that, in combination with the stereotypes, obtain the lowest perplexity score, while in Table 8, we report the top three most prevalent shared stereotypes across identities within each category for each model.
Figure 5 illustrates the logarithm of normalized perplexity scores across the four categories – religion, gender, nationality, and disability – indicating the distribution of these scores for the analyzed language models.
Figure 6 reports the perplexity-based filtering of SoFa stereotypes. The filtering is based on a threshold, specifically averaging perplexity scores from each model and creating a histogram to retain only stereotypes in selected bins exhibiting the lowest scores. Since our evaluation is based on identifying and quantifying perplexity peaks (both low and high), keeping statements in the dataset that are unlikely according to the models would offset the results. Other works in the literature also perform discarding statements with high perplexity scores to remove unlikely instances, e.g., Barikeri et al. (2021).
Model Religion Id. Gender Id. Nationality Id. Disability Id.
BLOOM-560m Jews, Taoists Women, Men Chinese, Japanese Normal people, Deafs, Disableds
BLOOM-3b Protestants, Priests, Taoists Men, Women, Trans guys Chinese, Japanese, Ecuadorians Little people, Disableds, Slow learners
GPT2-base Jews, Muslims, Orthodoxes Brotherboys, Trans males, Trans females Britishes, Beninese, Swatis Slow learners, Normal people, Lame people
GPT2-medium Jews, Muslims, Orthodoxes Brotherboys, Bi people, Enbies Britishes, Americans, Lesothans Slow learners, Midgets, Deviant people
XLNET-base-cased Catholics, Bhikkhus, Orthodoxes Men, Males, Bigender people Vietnamese, Sudanese, Israelis Short-statured people, Wheelchair users, People with depression
XLNET-large-cased Buddhist monks Males, Brotherboys, Enbies Mauritanians, Congolese, Japanese Invalids, Quadriplegics, Mongoloids
BART-base Bhikkhus, Mormons, Muslims Men, Trans females, Transgender females Nigeriens, Jamaicans, Serbians Differently abled people, Slow learners, Mongoloids
BART-large Mormons, Jews, Catholics Men, Trans men, Trans people Hispanic people, Iranians, Russians Normal people, Wheelchair users, Slow learners
LLAMA2-7b Jews, Christians Trans males, Women, Bi people Indians, Spanishes, Irishes Little people, Invalids, Normal people
LLAMA2-13b Christians, Jews Trans males, Trans females, Trans guys Indians, Turkishes, Portuguese Invalids, Little people, Retarded people
Table 7: Identities most frequently associated with stereotypes by category.
Model Religion St. Gender St. Nationality St. Disability St.
BLOOM-560m burned complain burned thrash around
suck suck loot aren’t athletes
aren’t ethical aren’t handy unemployed are handicap
BLOOM-3b burned complain unemployed aren’t athletes
screetch denied freedoms loot can’t jump
suck aren’t valid complain aren’t serious
GPT2-base homophobic complain burned are unlucky
suck suck unemployed making jokes
burned deserve less suck are shameful
GPT2-medium homophobic complain unemployed are orphans
suck deserve less loot are barbaric
burned are ineffective suck making jokes
XLNET-base steal gold are helpless are misinformed should be hurt
love to rob people scream when raped are hated by cops are suicide bombers
screetch exist only for sex were once enslaved are skinny
XLNET-large shoot innocent people like to flaunt genetically enjoy certain things like to kill people
rip people off are like animals kidnap people aren’t important
live in a crappy country are fools cheat the system always kill kids
BART-base come from horrible countries should keep their opinions to themselves stopped myself from rape are like skeletons
think are the best race bring up gender for no reason are pushy are dumb for cutting themselves
will chase you for what you have think they’re smart run fast when in trouble spitting on a disable person is ok
BART-large homophobic complain suck are contagious
burned behave badly unemployed behave badly
are poisonous hold no value burned deserve nothing
LLAMA2-7b died young accomplish nothing bother people hurt others
accomplish nothing hate them was killed indicate inferiority
marry young girls cannot help loot submit to anything
LLAMA2-13b died young accomplish nothing bother people hurt others
look albino hurt themselves hurt people are bad
lost wwii rule over everything hurt them behave badly
Table 8: Top-3 strongest stereotypes by category, i.e., the ones obtaining lowest DDS according to Equation 5.
Figure 3: Probes for the stereotype: [Identity] are all terrorists, where the different identities belonging to the category in consideration, in this case religion, are reported on the x-axis.
Figure 4: Stacked SoFa scores by category.
(a) Religion
(b) Gender
(c) Nationality
(d) Disability
Figure 5: Violin plots of $Log_{10}(PPL*)$ by category.
(a) Starting histogram.
(b) Resulting histogram after cutting at a threshold of $150$.
Figure 6: Perplexity-based filtering of SoFa stereotypes.